# Towards Interpretable and Efficient Attention: Compressing All by Contracting a Few

### v2

@@ 4.2 Evaluations on real-world visual tasks
We pretrain CBT models on the ImageNet-1K dataset, and finetune them on several downstream datasets. The top-1 accuracy on the validation sets is reported in Table 3. In particular, our CBT-Small achieves comparable top-1 accuracy to ViT-S using only 30% of the parameters and 40% of the FLOPs. Compared to CRATE models built upon [16], our CBT models (with convolutional embedding layers) perform remarkably better while using fewer parameters and FLOPs.
We also conduct fair comparisons across different attention mechanisms, which can be regarded as variants of CBSA, and report the results as in Table 4. In this setting, our CBT models remain competitive with CRATE while computing significantly fewer pairwise similarities. These results confirm that, from MSSA to CBSA and then to TSSA, the number of pairwise similarity computations decreases at the cost of some performance. Throughput comparisons on high-resolution images ($512\times 512$) are presented in Table 7, showing that our method consistently achieves superior training and inference efficiency.
To investigate the potential of applying CBSA to large-scale pretraining, we preliminarily finetune ViT models pretrained on ImageNet-21K [^14] The checkpoint is obtained from the timm [42] library. by adapting their attention blocks into the CBSA style. For comparison, we also adapt them into linear attention, with results shown in Fig. 9. Although CBSA deviates more from MHSA than linear attention and is consequently harder to adapt from pretrained ViTs, CBSA achieves performance comparable to linear attention with nearly the same FLOPs. Moreover, when adapted to Agent Attention, CBSA surpasses linear attention, which has been carefully experimented in [30].
For semantic segmentation, we build CBT decoders by stacking CBSA layers without feed-forward modules on top of ViT encoders, following the design of [31]. We evaluate them on the ADE20K dataset [43], with results shown in the left panel of Fig. 8. Our CBT decoder consistently surpasses both white-box (DEPICT) and black-box (Segmenter) counterparts that rely on softmax attention. In particular, the best-performing CBT decoder improves upon Segmenter by $1.5\%$ mIoU while using only $20\%$ of the FLOPs and $0.06\%$ of the pairwise similarities in the decoder.
@@ 5 Related work
Efficient attention mechanisms can be roughly divided into two categories: sparse attention and linear attention [44]. Sparse attention approaches sparsify the attention matrix proactively by restricting the attention span to random, fixed [45, 19], learnable [46, 47] patterns, or their combinations. Linear attention [20, 48] decomposes the attention matrix into a product of two low-rank matrices and thus avoids its explicit computation via the associative property of multiplication. The idea of using representative tokens has been applied to both. Global tokens or memory can be introduced in sparse attention to maintain global connectivity, thereby further shrinking the attention span [49, 50]. Meanwhile, various works are implicitly sparse for reducing the number of input tokens [51, 52, 53, 54]. Agent tokens [30] and landmarks [28] can also be incorporated into linear attention from different perspectives.
Our CBSA distinguishes itself from previous efficient attention mechanisms by being inherently interpretable and derived from an optimization objective that efficiently compresses input tokens toward low-dimensional structures. Moreover, it can not only be viewed as sparse attention or linear attention through the representatives, but also mathematically generalizes linear attention and channel attention as special cases.
@@ 6 Conclusion
We proposed an optimization objective that compresses all input tokens by contracting a small number of representatives, thereby unifying the investigation towards interpretability and efficiency. By unrolling the gradient optimization steps for this objective, we derived an inherently interpretable and efficient attention mechanism, called CBSA. We found that many existing attention mechanisms can be generalized within the CBSA framework by specifying the number and structure of representatives, thus revealing their fundamental connections. We further validated the effectiveness of CBSA through extensive experiments on visual tasks. Therefore, we demonstrated that the preliminary framework established in this work offers a promising direction for exploring new attention mechanisms in an interpretable and efficient way.
@@ Appendix A Representative initialization
Besides pooling the input tokens, another intuitive choice for representative initialization is to set them as learnable (i.e., trainable) parameters, similar to the object queries in [32], the class embeddings in [40], and the registers in [55]. However, in this case we found that the attention matrix in the extraction (cross-attention) step [^15] That is, $\operatorname{rank}\left(\operatorname{softmax}\left((\boldsymbol{U}_{k}^{\top}\boldsymbol{Z})^{\top}(\boldsymbol{U}_{k}\boldsymbol{Q})\right)\right)$. is extremely low rank. For example, although its maximal rank can be $m=p=\frac{d}{H}=64$ in CBT models, it typically remains around $5$. This indicates that far fewer initialized representatives are being utilized than expected, leading to unsatisfactory performance of CBSA.
@@ Appendix B Derivations
$=$ $\sum_{k=1}^{K}\boldsymbol{U}_{k}F\left(\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k}^{2}\boldsymbol{L}_{k}^{\top}\right)\boldsymbol{U}_{k}^{\top}\boldsymbol{Z}$ (19)
$=$ $\sum_{k=1}^{K}\boldsymbol{U}_{k}F\left((\boldsymbol{U}_{k}^{\top}\boldsymbol{Z})(\boldsymbol{U}_{k}^{\top}\boldsymbol{Z})^{\top}\right)\boldsymbol{U}_{k}^{\top}\boldsymbol{Z}$ (20)
@@ Appendix C More experimental results
Training setup. We train the CBT models in Table 3 and all models in Table 4 150 epochs with the Lion optimizer [57]. The learning rate is $2.0\times 10^{-4}$, the weight decay coefficient is $0.05$, and the batch size is $256$. We also incorporate a warm-up strategy over the first $20$ epochs. For data augmentation, we adopt a rather simple choice: just random cropping and random horizontal flipping. We apply label smoothing with a smoothing coefficient of $0.1$. For fine-tuning, we use the AdamW optimizer [58], a learning rate of $5\times 10^{-5}$, weight decay of 0.010.0-1, and batch size $64$. The settings above are largely inherited from [16].
Quantitative segmentation properties. Following prior works [7, 41], we evaluate the zero-shot segmentation performance of ImageNet-1K pretrained models on the PASCAL VOC12 validation set [59]. Specifically, we assess the best-performing attention maps of the [CLS] token, but in a simplified setting that considers only three segmentation targets, instead of fine-grained semantic classes. In Table 5, we find that the hybrid model has consistently better segmentation performance. We also measure the segmentation performance of different layers in Table 6.
@@
Table 7: Throughput comparisons. The results below are experimented on the CIFAR-10 dataset with an image resolution of $512\times 512$.
im/sec CBT-T ViT-T CRATE-T ToST-T CBT-S ViT-S CRATE-S ToST-S
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Gram matrices of basis vectors modeled by the layers of CBT-S. According to our assumption in Section 2, these matrices should approximate identity matrices.
@@ Appendix D PyTorch implementation
1
1class CBSA($n$.Module):
2 def __init__(self, dim, heads, dim_head):
3 super().__init__()
4 inner_dim = heads * dim_head
5 self.heads = heads
6 self.dim_head = dim_head
7 self.scale = dim_head ** -0.5
8 self.attend = $n$.Softmax(dim=-1)
9 # subspace bases
10 self.proj = $n$.Parameter(self.scale * torch.randn(inner_dim, dim))
11 # over-parameterization
12 self.to_out = $n$.Linear(inner_dim, dim)
13 # step-sizes
14 self.ss_x = $n$.Parameter(torch.randn(heads, 1, 1))
15 self.ss_rep = $n$.Parameter(torch.randn(heads, 1, 1))
16
17 def attention(self, query, key, value):
18 dots = (query @ key.transpose(-1, -2)) * self.scale
19 attn = self.attend(dots)
20 out = attn @ value
21 return out, attn
22
23 def forward(self, x):
24 b, n, c = x.shape
25 # normalize the basis vectors
26 proj = self.proj / self.proj.norm(dim=-1, keepdim=True)
27 # projection onto subspaces
28 w = rearrange(x @ proj.t(), ’b n (h k) -> b h n k’, h=self.heads)
29 # representative initialization (couple them with the basis vectors but preserved the norm)
30 rep = rearrange(self.proj, ’(h k) d -> h k d’, h=self.heads)
31 rep = rep @ rearrange(proj, ’(h k) d -> h k d’, h=self.heads).transpose(-1, -2)
32 rep = rep.expand(b, -1, -1, -1)
33 # representative extraction
34 rep_delta, attn = self.attention(rep, w, w)
35 rep = rep + self.ss_rep * rep_delta
36 # representative contraction
37 x_delta, _ = self.attention(rep, rep, rep)
38 # contraction broadcast
39 x_delta = attn.transpose(-1, -2) @ x_delta
40 x_delta = self.ss_x * x_delta
41 x_delta = rearrange(x_delta, ’b h n k -> b n (h k)’)
42 # projection back to the ambient space
43 return self.to_out(x_delta)
Algorithm 2 PyTorch implementation of CBSA in Fig. 8
1
1class CBSA($n$.Module):
2 def __init__(
3 self,
4 dim: int,
5 num_heads: int = 8,
6 qkv_bias: bool = False,
7 proj_bias: bool = True,
8 ) -> None:
9 super().__init__()
10 assert dim % num_heads == 0, ’dim should be divisible by num_heads’
11 self.num_heads = num_heads
12 self.head_dim = dim // num_heads
13 self.scale = self.head_dim ** -0.5
14 self.pool = $n$.AdaptiveAvgPool2d(output_size=(8, 8))
15 self.qkv = $n$.Linear(dim, dim * 3, bias=qkv_bias)
16 self.proj = $n$.Linear(dim, dim, bias=proj_bias)
17
18 def attention(self, query, key, value):
19 query = query * self.scale
20 attn = query @ key.transpose(-2, -1)
21 attn = attn.softmax(dim=-1)
22 out = attn @ value
23 return out, attn
24
25 def forward(self, X: torch.Tensor) -> torch.Tensor:
26 B, N, C = X.shape
27 X_qkv = self.qkv(X).reshape(B, N, 3, C).permute(2, 0, 1, 3)
28 X_q, X_k, X_v = X_qkv.unbind(0)
29 H = W = int(N ** 0.5)
30 Q = self.pool(X_q[:, :-1, :].reshape(B, H, W, C).permute(0, 3, 1, 2)).reshape(B, C, -1).permute(0, 2, 1)
31 X_q = X_q.reshape(B, N, self.num_heads, self.head_dim).permute(0, 2, 1, 3)
32 X_k = X_k.reshape(B, N, self.num_heads, self.head_dim).permute(0, 2, 1, 3)
33 X_v = X_v.reshape(B, N, self.num_heads, self.head_dim).permute(0, 2, 1, 3)
34 Q = Q.reshape(B, 64, self.num_heads, self.head_dim).permute(0, 2, 1, 3)
35 Q_delta, _ = self.attention(Q, X_k, X_v)
36 Q = Q + Q_delta
37 X_delta, _ = self.attention(Q, Q, Q)
38 attn = ((X_q * self.scale) @ Q.transpose(-1, -2)).softmax(dim=-1)
39 X = attn @ X_delta
40 X = X.transpose(1, 2).reshape(B, N, C)
41 X = self.proj(X)
42 return X
Algorithm 3 PyTorch implementation of CBSA in Fig. 9

### v3

@@ 4.2 Evaluations on real-world visual tasks
We pretrain CBT models on the ImageNet-1k dataset, and finetune them on several downstream datasets. The top-1 accuracy on validation sets is reported in Table 3. In particular, our CBT-Small achieves comparable top-1 accuracy to ViT-S using only 30% of the parameters and 40% of the FLOPs. Compared to CRATE models, our CBTs (with convolutional embedding layers) perform remarkably better while using fewer parameters and FLOPs.
We also conduct a set of fair comparisons across different attention mechanisms, which can be regarded as variants of CBSA, and report the results in Table 4. In this setting, our CBT models remain competitive with CRATE while computing significantly fewer pairwise similarities. These results confirm that, from MSSA to CBSA and then to TSSA, the number of pairwise similarity computations decreases at the cost of performance sacrifice. In addition, we present the throughput comparisons on high-resolution images (e.g., $512\times 512$) in Table 7 (Appendix C), showing that our methods consistently achieves superior training and inference efficiency.
To investigate the potential of applying CBSA to large-scale pretraining, we preliminarily finetune ViT models pretrained on ImageNet-21k [^13] The checkpoint is obtained from the timm [53] library. by adapting their attention blocks into the CBSA style. For comparison, we also adapt them into linear attention. Experimental results are shown in Fig. 9. Although CBSA deviates more from MHSA than linear attention and is consequently harder to adapt from pretrained ViTs, CBSA achieves comparable performance to that of linear attention with nearly the same FLOPs. Interestingly, when the contraction step is removed, CBSA surpasses linear attention, which has been extensively investigated in [39].
For semantic segmentation, following the design of DEPICT [36], we build CBT decoders by stacking CBSA layers without the feed-forward modules on the top of ViT encoders. We evaluate the performance of them on the ADE20K dataset [54] and show the results in the left panel of Fig. 8. Clearly, our CBT decoder consistently surpasses both white-box (DEPICT) and black-box (Segmenter) counterparts that rely on softmax attention. In particular, the best-performing CBT decoder improves upon Segmenter by $1.5\%$ mIoU while using merely $20\%$ of the FLOPs and $0.06\%$ of the pairwise similarities in the decoder.
@@ 5 Related work
Efficient attention mechanisms can be roughly divided into two categories: sparse attention and linear attention [55]. Approaches in sparse attention sparsify the attention matrix proactively by restricting the attention span to either random, or fixed [56, 20], or learnable [57, 58] patterns, or their combinations. Approaches in linear attention [21, 59] decompose the attention matrix into a product of two low-rank matrices and thus avoids its explicit computation via the associative property of multiplication. The idea of using representative tokens has been applied to both. Global tokens or memory can be introduced in sparse attention to maintain global connectivity, thereby further shrinking the attention span [60, 61]. Meanwhile, Agent tokens [39] and landmarks [32] can also be incorporated into linear attention from different perspectives.
Our CBSA distinguishes itself from previous efficient attention mechanisms by being inherently interpretable and derived from an optimization objective which efficiently compresses input tokens towards low-dimensional structures. Moreover, it can not only be viewed as sparse attention or linear attention for leveraging the representatives, but also mathematically generalizes softmax attention, linear attention, and channel attention as its special cases.
@@ 6 Conclusion
We have proposed an optimization objective for deriving attention mechanism and unifying the investigation towards the interpretability and the efficiency. By unrolling the gradient optimization steps of this objective, we derived an inherently interpretable and efficient attention mechanism, called Contract-and-Broadcast Self-Attention (CBSA). We found that our CBSA covers the instantiations of softmax attention, linear attention, and channel attention by changing the number and structure of representatives, thus revealing their fundamental connections. We validated the effectiveness of our CBSA through extensive experiments on visual tasks. We believe that the preliminary framework established in this work offers a promising direction for exploring a unified formula for existing attention mechanisms as well as new attention mechanisms in an inherently interpretable way.
Acknowledgments and Disclosure of Funding
This work is supported by the National Natural Science Foundation of China under Grant Nos. 62576048 and 61876022. C.-G. Li is the corresponding author.
@@ Appendix A Representative initialization
Besides pooling the input tokens, another intuitive choice for representative initialization is to set them as learnable (i.e., trainable) parameters, similar to the object queries in [62], the class embeddings in [51], and the registers in [63]. However, in this case we found that the attention matrix in the extraction (cross-attention) step [^14] That is, $\operatorname{rank}\left(\operatorname{softmax}\left((\boldsymbol{U}_{k}^{\top}\boldsymbol{Z})^{\top}(\boldsymbol{U}_{k}\boldsymbol{Q})\right)\right)$. is low rank. For example, although its maximal rank can be $m=p=\frac{d}{H}=64$ in CBT models, it typically remains around $5$. This indicates that far fewer initialized representatives are being utilized than expected, leading to unsatisfactory performance of CBSA.
@@ Appendix B Derivations
$=$ $\sum_{k=1}^{K}\boldsymbol{U}_{k}\mathcal{F}\left(\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k}^{2}\boldsymbol{L}_{k}^{\top}\right)\boldsymbol{U}_{k}^{\top}\boldsymbol{Z}$ (19)
$=$ $\sum_{k=1}^{K}\boldsymbol{U}_{k}\mathcal{F}\left((\boldsymbol{U}_{k}^{\top}\boldsymbol{Z})(\boldsymbol{U}_{k}^{\top}\boldsymbol{Z})^{\top}\right)\boldsymbol{U}_{k}^{\top}\boldsymbol{Z}.$ (20)
Note that an arbitrary set of orthogonal representatives can be expressed as the principal directions under an orthogonal rotation and scaling, i.e.,
$\boldsymbol{U}_{k}^{\top}\boldsymbol{Q}=\boldsymbol{P}_{k}\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k}\boldsymbol{\Lambda}_{k}=\boldsymbol{P}_{k}\boldsymbol{U}_{k}^{\top}\boldsymbol{Z}\boldsymbol{R}_{k}\boldsymbol{\Lambda}_{k},$ (21)
where $\boldsymbol{P}_{k}\in\mathcal{O}{(p)}$ denotes an orthogonal rotation within the $k$-th subspace, and $\boldsymbol{\Lambda}_{k}$ is a diagonal matrix of size $p$. Generalizing the way to form the representatives in (4) from $\boldsymbol{U}_{k}^{\top}\boldsymbol{Q}=\boldsymbol{U}_{k}^{\top}\boldsymbol{Z}\boldsymbol{A}_{k}$ to $\boldsymbol{U}_{k}^{\top}\boldsymbol{Q}=\boldsymbol{B}_{k}\boldsymbol{U}_{k}^{\top}\boldsymbol{Z}\boldsymbol{A}_{k}$, where $\boldsymbol{B}_{k}\in\mathbb{R}^{d\times d}$, we have:
$\nabla_{\boldsymbol{Z}}R_{c}(\boldsymbol{Q}\mid\boldsymbol{U}_{[K]})\propto\sum_{k=1}^{K}\boldsymbol{U}_{k}\boldsymbol{B}_{k}^{\top}\boldsymbol{U}_{k}^{\top}\boldsymbol{Q}\left(\mathbf{I}_{m}+\frac{p}{m\epsilon^{2}}(\boldsymbol{U}_{k}^{\top}\boldsymbol{Q})^{\top}(\boldsymbol{U}_{k}^{\top}\boldsymbol{Q})\right)^{-1}\boldsymbol{A}_{k}^{\top}.$ (22)
Similarly, by substituting $\boldsymbol{U}_{k}^{\top}\boldsymbol{Q}=\boldsymbol{P}_{k}\boldsymbol{L}_{k}\boldsymbol{\Sigma}_{k}\boldsymbol{\Lambda}_{k}$, $\boldsymbol{A}_{k}=\boldsymbol{R}_{k}\boldsymbol{\Lambda}_{k}$ and $\boldsymbol{B}_{k}=\boldsymbol{P}_{k}$ into (22), we obtain
$\sum_{k=1}^{K}\boldsymbol{U}_{k}\boldsymbol{L}_{k}\boldsymbol{\Lambda}_{k}\left(\mathbf{I}_{m}+\frac{1}{\epsilon^{2}}\boldsymbol{\Lambda}_{k}^{2}\boldsymbol{\Sigma}_{k}^{2}\right)^{-1}\boldsymbol{\Lambda}_{k}\boldsymbol{L}_{k}^{\top}\boldsymbol{U}_{k}^{\top}\boldsymbol{Z}.$ (23)
@@ Appendix C More experimental results
Training setup. We train the CBT models in Table 3 and all models in Table 4 150 epochs with the Lion optimizer [64]. The learning rate is $2.0\times 10^{-4}$, the weight decay coefficient is $0.05$, and the batch size is $256$. We also incorporate a warm-up strategy over the first $20$ epochs. For data augmentation, we adopt a rather simple choice: just random cropping and random horizontal flipping. We apply label smoothing with a smoothing coefficient of $0.1$. For fine-tuning, we use the AdamW optimizer [65], a learning rate of $5\times 10^{-5}$, weight decay of $0.01$, and batch size $64$. The settings above are largely inherited from [23].
Quantitative segmentation properties. Following prior works [6, 52], we evaluate the zero-shot segmentation performance of ImageNet-1K pretrained models on the PASCAL VOC12 validation set [66]. Specifically, we assess the best-performing attention maps of the [CLS] token, but in a simplified setting that considers only three segmentation targets, instead of fine-grained semantic classes. In Table 5, we find that the hybrid model has consistently better segmentation performance. We also measure the segmentation performance of different layers and report the results in Table 6.
@@
Table 7: Throughput comparisons. The results below are obtained from experiments on dataset CIFAR-10 with an image resolution of $512\times 512$.
Images / sec CBT-T ViT-T CRATE-T ToST-T CBT-S ViT-S CRATE-S ToST-S